### v2

@@ 2.2 HIPPO Achieves Robust Performance in Intra-species PPI prediction
Figure 3: Protein embeddings and attention analysis for PPI prediction. (A) The left-most column highlights residue-level embeddings colored by amino acid biophysical properties. The middle column annotates embeddings according to protein domain classifications. The right-most column distinguishes proteins by lineage or organism classifications. The comparative visualization clearly illustrates improved clustering and distinct separation patterns in the lower row (w/ Hierarchy), demonstrating the enhanced representational capability achieved by incorporating hierarchical constraints during training. t-SNE projections visualize information extracted by protein representation models (upper row: without hierarchical pre-training on Swiss-Prot, and lower row: with hierarchical pre-training on Swiss-Prot). Residue: Visualization of residue embedding. Two-dimensional projection of amino acid token embeddings colored and shaped by physicochemical properties (e.g., polarity, hydrophobicity, and size). The separation among groups reflects the model’s ability to learn property-aware representations. Domain: Visualization of 4 different domain embedding( Protein kinase domain, WD40 repeats, Immunoglobulin domain, Ankyrin repeat). Lineage: Visualization of the protein globel embeddings on the Beta-lactamase family PF00144 from Pfam. (C) Identification of binding site(PDB id: 3VZ9). The area in red represents the predicted binding site region, while the blue area represents the missed area of binding sites. The white area depicts the non-essential area of the protein. (D)Distributions of AUPR scores on 5 PPI types. Bar plots showing F1 scores of four models (Ours, ESM-2, GNN-PPI, and MASSA) across five distinct interaction types (Reaction, Binding, PTMod, Activation, and Catalysis) under DFS, BFS, and random sampling strategies on the SHS27k dataset.
@@ 2.4 HIPPO Extends to Cross-Species PPI Prediction with Leading Performance
Cross-species protein–protein interaction (PPI) prediction addresses the challenge of inferring interactions in less-studied species by leveraging knowledge learned from well-studied organisms, such as humans. Unlike traditional approaches that rely on species-specific training data, our framework transfer predictive power across species boundaries, enabling PPI prediction where experimental data are scarce or unavailable. In this section, we (1) systematically evaluate the model’s ability to predict PPIs in 5 species after training on human dataset, (2) benchmark against existing methods, and (3) analyze the factors underlying successful cross-species generalization.
As summarized in Fig. 4, our model consistently outperforms existing methods, including GNN-PPI, ProteinBERT, ESM-2, PIPR, and INTREPPID, across all cross-species datasets. The proposed approach achieves the highest F1 scores in each species, demonstrating its robustness and superior predictive power in challenging cross-species settings. These results highlight the capacity of our hierarchical, multimodal framework to learn transferable protein representations that generalize well to novel proteomes.
The strong cross-species performance suggests that the model effectively captures conserved functional features and interaction patterns shared among diverse organisms. This broad applicability is particularly valuable for predicting PPIs in non-model species, where experimental data are limited. Overall, these findings underscore the potential of our approach for large-scale, cross-species interactome mapping.
Cross-species protein–protein interaction (PPI) prediction addresses the challenge of inferring interactions in less-studied species by leveraging knowledge learned from well-studied organisms, such as humans. Unlike traditional approaches that rely on species-specific training data, our framework transfers predictive power across species boundaries, enabling PPI prediction where experimental data are scarce or unavailable. In this section, we (1) systematically evaluate the model’s ability to predict PPIs in six species after training on human datasets-specifically, Escherichia coli, Saccharomyces cerevisiae (yeast), Mus musculus (mouse), Caenorhabditis elegans, Arabidopsis thaliana, and Drosophila melanogaster, all of which represent widely studied model organisms in genetics, developmental biology, and plant science; (2) benchmark against existing methods; and (3) analyze the factors underlying successful cross-species generalization.
As summarized in Fig. 4, our model consistently outperforms existing methods, including GNN-PPI, ProteinBERT, ESM-2, PIPR, and INTREPPID, across all cross-species datasets. The proposed approach achieves the highest F1 scores in each species, demonstrating its robustness and superior predictive power in challenging cross-species settings. These results highlight the capacity of our hierarchical, multimodal framework to learn transferable protein representations that generalize well to novel proteomes. The performance suggests that the model effectively captures conserved functional features and interaction patterns shared among diverse organisms. Notably, the model’s robust results in intra-species PPI prediction provide a solid foundation for its success in cross-species scenarios, as both settings benefit from the hierarchical and transferable representations learned by the framework. This broad applicability is particularly valuable for predicting PPIs in non-model species, where experimental data are limited. Overall, these findings underscore the potential of our approach for large-scale, cross-species interactome mapping.

### v3

@@ 2.2 HIPPO Achieves Robust Performance in Intra-species PPI prediction
Figure 3: Protein embeddings and attention analysis for PPI prediction. (A) The left column highlights residue-level embeddings colored by amino acid biophysical properties. The middle column annotates embeddings according to protein domain classifications. The right column distinguishes proteins by lineage or organism classifications. The comparative visualization clearly illustrates improved clustering and distinct separation patterns in the lower row (w/ Hierarchy). t-SNE projections visualize information extracted by protein representation models (upper row: without hierarchical pre-training on Swiss-Prot, and lower row: with hierarchical pre-training on Swiss-Prot). Residue: Visualization of residue embedding. Two-dimensional projection of amino acid token embeddings colored and shaped by physicochemical properties (e.g., polarity, hydrophobicity, and size). The separation among groups reflects the model’s ability to learn property-aware representations. Domain: Visualization of 4 different domain embedding( Protein kinase domain, WD40 repeats, Immunoglobulin domain, Ankyrin repeat). Lineage: Visualization of three different lineage embeddings (Actinobacteria, Bacteroidetes, and Firmicutes) on the Beta-lactamase family PF00144 from Pfam. (C) Identification of binding site(PDB id: 3VZ9). The area in red represents the predicted binding site region, while the blue area represents the missed area of binding sites. The white area depicts the non-essential area of the protein. (D)Distributions of AUPR scores on 5 PPI types. Bar plots showing F1 scores of four models (Ours, ESM-2, GNN-PPI, and MASSA) across five distinct interaction types (Reaction, Binding, PTMod, Activation, and Catalysis) under DFS sampling strategies on the SHS27k dataset. (E) Bar plots display the overlap rates for binding site prediction across 32 protein complexes under three model configurations: hierarchical information combined with flat annotations (+/++/++ / +), hierarchical information without flat annotations (+/+/-+ / -), and neither hierarchical nor flat annotations (/-/-- / -).
@@ 2.4 HIPPO Extends to Cross-Species PPI Prediction with Leading Performance
Cross-species protein-protein interaction (PPI) prediction addresses the challenge of inferring interactions in less-studied species by leveraging knowledge learned from well-studied organisms, such as humans. Unlike traditional approaches that rely on species-specific training data, our framework transfers predictive power across species boundaries, enabling PPI prediction where experimental data are scarce or unavailable. In this section, we (1) systematically evaluate the model’s ability to predict PPIs in six species after training on human datasets-specifically, Escherichia coli, Saccharomyces cerevisiae (yeast), Mus musculus (mouse), Caenorhabditis elegans, Arabidopsis thaliana, and Drosophila melanogaster, all of which represent widely studied model organisms in genetics, developmental biology, and plant science; (2) benchmark against existing methods; and (3) analyze the factors underlying successful cross-species generalization.
As summarized in Fig. 4, our model consistently outperforms existing methods, including GNN-PPI, ProteinBERT, ESM-2, PIPR, and INTREPPID, across all cross-species datasets. The proposed approach achieves the highest F1 scores in each species, demonstrating its robustness and superior predictive power in challenging cross-species settings. These results highlight the capacity of our hierarchical, multimodal framework to learn transferable protein representations that generalize well to novel proteomes. The performance suggests that the model effectively captures conserved functional features and interaction patterns shared among diverse organisms. Notably, the model’s robust results in intra-species PPI prediction provide a solid foundation for its success in cross-species scenarios, as both settings benefit from the hierarchical and transferable representations learned by HIPPO. This broad applicability is particularly valuable for predicting PPIs in non-model species, where experimental data are limited. Overall, these findings underscore the potential of our approach for large-scale, cross-species interactome mapping.